### v1

# Interval Decomposition of Infinite Persistence Modules over a Principal Ideal Domain

Abstract.
We study pointwise free and finitely-generated persistence modules over a principal ideal domain, indexed by a (possibly infinite) totally-ordered poset category. We show that such persistence modules admit interval decompositions if and only if every structure map has free cokernel. We also show that, in torsion-free settings, the integer persistent homology module of a filtration of topological spaces admits an interval decomposition if and only if the associated persistence diagram is invariant to the choice of coefficient field. These results generalize prior work where the indexing category is finite.
2020 Mathematics Subject Classification:
Primary MSC 62R40, MSC 55N31; Secondary MSC 55-08
1. Introduction
Let $R$ be a principal ideal domain (PID). A persistence module with coefficients in $R$ is a functor $f:I\to R\mathrm{-Mod}$, where the indexing category $I$ is a totally-ordered poset category. Persistence modules arise from and have been widely studied in persistent homology, in which setting $R$ is typically a field and $f$ maps every index $i\in I$ to a finite-dimensional vector space over $R$. In this setting, any persistence module admits an interval decomposition; that is, it can be decomposed as a direct sum of indecomposables known as “interval modules”. Interval decompositions are key to persistent homology, in that they capture important topological structure within data. If one removes the condition that the coefficient ring $R$ be a field, then interval decompositions are no longer guaranteed to exist.
In this paper, we study persistence modules that are pointwise free and finitely-generated, in which the coefficient ring is a PID and the indexing category is any totally-ordered poset category. We introduce necessary and sufficient conditions for such modules to decompose (see Section 2 for definitions):
Theorem 1.1 (Main result).
Let $f:I\to R\mathrm{-Mod}$ be a pointwise free and finitely-generated persistence module, where $I$ may be infinite and $R$ is a PID. Then $f$ admits an interval decomposition if and only if every map $f(a\leq b)$ has free cokernel.
The existence of interval decompositions for modules with field coefficients has been widely attributed to Gabriel [6], whose work showed that indecomposable representations of quivers of type $A_{n}$ have only finitely many isomorphism classes, which coincide exactly with interval modules. This result has since been generalized to persistence modules indexed by any totally-ordered set [4], as well as a small category [2]. Interval decompositions have also been generalized to zigzag persistence. Carlson and de Silva [3] showed that interval decompositions exist for finitely-indexed zigzag persistence modules; Botnan [1] later showed the same in the infinitely-indexed setting. Hang and Mio [9] generalized interval decomposition to tame correspondence modules.
The problem of decomposing modules with PID coefficients is motivated by growing mathematical and scientific interest in modules with non-field coefficients, such as in Conley index theory [11, 10], generalized persistence diagrams [14], projective coordinatization [5, 15], and learned representations in neural networks [17].
However, this problem presents nontrivial structural challenges which are absent from the works described above. First, the endomorphism rings of indecomposable modules with PID coefficients are nonlocal, which complicates the approach of [2]. In particular, Theorem 1.1 and Lemma 2.1 of that work bear heavily on the proof of interval decomposition [2, Theorem 1.2], but all three results fail over a PID. Second, arbitrary submodules of free $R$-modules need not have complements, which complicates the approach of [4]. For example, the limiting object $V_{I}^{-}$ lacks a complement to $V_{I}^{+}$ in general, which impedes the construction in [4, Section 5], and creates additional challenges in [4, Section 6].
To address these problems, we abstract certain aspects of the algebraic structure, focusing instead on the structure of subobject lattices. Specifically, we show that the functorial properties of the saecular lattice [10, 8] can be used to build decompositions constructively from images and kernels.
Our work complements and builds on [10], which established necessary and sufficient conditions for interval decompositions of persistence modules over PIDs with finite indexing categories. The core element of [10] is a polynomial-time algorithm to compute decompositions explicitly. While key ideas in [10] remain relevant, the generality of our indexing category requires nontrivial adaptation.
2. Background and notation
In this section, we present the necessary background for persistence modules and set up the notation of this paper. We also provide a brief discussion on finitely-indexed persistence modules, which will be relevant to prove the main result, Theorem 1.1. See [7, 16, 13] for more on persistence modules and homology.
2.1. Persistence modules
Let $R$ be a ring and $I$ a totally-ordered poset category.
Definition 2.1.
A persistence module indexed by $I$ with coefficients in $R$ is a functor $f:I\to R\mathrm{-Mod}$ from $I$ to the category of $R$-modules.
In other words, $f$ consists of (1) a family $\{f_{i}\}_{i\in I}$ of $R$-modules and (2) morphisms (which we call structure maps) $f(i\leq j):f_{i}\to f_{j}$ whenever $i\leq j$, in which $f(i\leq i)=\mathrm{id}_{f_{i}}$ for all $i\in I$ and $f(i\leq k)=f(j\leq k)\circ f(i\leq j)$ whenever $i\leq j\leq k$.
For a persistence module $f:I\to R\mathrm{-Mod}$, we denote $f_{a}$ as $f$ evaluated at $a$, and $f(a\leq b):f_{a}\to f_{b}$ as the structure map between $f_{a}$ and $f_{b}$.
For the remainder of this work, unless otherwise stated, $f$ will denote a persistence module $f:I\to R\mathrm{-Mod}$ which is pointwise free and finitely-generated. That is, for each $i\in I$, we have that $f_{i}$ is a finitely-generated free $R$-module. We will assume, without loss of generality, that $I$ has a minimal and maximal element, which we denote -$\infty$ and $\infty$, respectively, and that $f_{-\infty}=f_{\infty}=0$. This assumption simplifies our analysis without loss of generality.
Definition 2.2.
Let $f,g:I\to R\mathrm{-Mod}$ be persistence modules. We say that $\phi:f\to g$ is a morphism between persistence modules if $\phi$ is a natural transformation. That is, $\phi$ consists of a family $\{\phi_{i}\}_{i\in I}$ of maps such that for all $a,b\in I$ where $a\leq b$, we have $\phi_{b}\circ f(a\leq b)=g(a\leq b)\circ\phi_{a}$. An isomorphism between persistence modules is a morphism with an inverse.
Definition 2.3.
Let $I$ be a totally-ordered poset. An interval is a subset $J\subseteq I$ such that for all $a,b\in J$ and $c\in I$ such that $a\leq c\leq b$, we have $c\in J$.
Definition 2.4.
An interval module is a persistence module of the form $h^{J}:I\to R\mathrm{-Mod}$, where $J$ is an interval,
[…]
An interval submodule of $f$ is any persistence submodule that is isomorphic to an interval module. An interval decomposition of $f$ is a direct-sum decomposition in which each summand is an interval submodule.
Remark 2.5.
$f$ admitting an interval decomposition is equivalent to the condition that, for all $a\leq b$, the matrix representation of $f(a\leq b)$ with respect to bases $\beta_{a}$ and $\beta_{b}$ is a matching matrix [^1] A matching matrix is a matrix with coefficients in $\{0,1\}$, where each row (respectively, each column) contains at most one nonzero entry.
Much of our discussion will focus on images and kernels, so we will use some convenient shorthand for these modules:
Definition 2.6.
Fix $a\in I$. For $x,y\in I$, we define
[…]
We also define
$\mathrm{IK}_{a}[x,y]=\mathrm{Im}[x,a]\cap\mathrm{Ker}[a,y].$
Remark 2.7.
For $y_{1},y_{2}\in I$ such that $y_{1}\leq y_{2}$, we have $\mathrm{Ker}[a,y_{1}]\subseteq\mathrm{Ker}[a,y_{2}]$. Similarly, for $x_{1},x_{2}\in I$ such that $x_{1}\leq x_{2}$, we have $\mathrm{Im}[x_{1},a]\subseteq\mathrm{Im}[x_{2},a]$.
Definition 2.8.
Let $M$ be a $R$-module, and let $A,B\subseteq M$ be submodules such that $A\subseteq B$. We say that $A$ has a complement in $B$ if we can find a submodule $C\subseteq B$ such that $A\oplus C=B$. We refer to $C$ as a complement of $A$ in $B$.
Lemma 2.9.
If $f$ is pointwise free and finitely-generated, then for a fixed $a\in I$, the family $\{\mathrm{Ker}[a,y]\}_{y\in I}$ of kernels out of $f_{a}$ is finite. Similarly, if the cokernels of structure maps are free, then the family $\{\mathrm{Im}[x,a]\}_{x\in I}$ of images into $f_{a}$ is finite.
We will use the following fact: if $M$ is an $R$-module, and $A\subseteq B\subseteq M$ where $A$ and $B$ are free, then $B/A$ is free if and only if $A$ has a complement in $B$.
Proof of Lemma 2.9.
We first show that the family $\{\mathrm{Ker}[a,y]\}_{y\in I}$ of kernels out of $f_{a}$ is finite. Recall that for $y\leq a$, we define $\mathrm{Ker}[a,y]=0$, while for $y>a$ we define $\mathrm{Ker}[a,y]=\ker(f(a\leq y))$. Let $y_{1},y_{2}\in I$ such that $a<y_{1}\leq y_{2}$. Then we have a short exact sequence
$0\to\mathrm{Ker}[a,y_{1}]\hookrightarrow\mathrm{Ker}[a,y_{2}]\xrightarrow{f(a\leq y_{1})}Z\to 0$
for some $Z\subseteq f_{y_{1}}$. Because $f_{y_{1}}$ is free, so too is $Z$. Therefore the sequence splits; in particular $\mathrm{Ker}[a,y_{1}]$ has a complement in $\mathrm{Ker}[a,y_{2}]$. If there exists an infinite sequence $\mathrm{Ker}[a,y_{1}]\subsetneq\mathrm{Ker}[a,y_{2}]\subsetneq\cdots$, then $f_{a}$ would therefore be infinite-rank — a contradiction. Thus, the family $\{\mathrm{Ker}[a,y]\}_{y\in I}$ of kernels out of $f_{a}$ is finite.
We now show that the family $\{\mathrm{Im}[x,a]\}_{x\in I}$ of images into $f_{a}$ is finite. Recall that for $x\geq a$, we define $\mathrm{Im}[x,a]=f_{a}$, while for $x<a$ we define $\mathrm{Im}[x,a]=\mathrm{Im}(f(x\leq a))$. Let $x_{1},x_{2}\in I$ be such that $x_{1}\leq x_{2}<a$. Then we have a short exact sequence
$0\to\mathrm{Im}[x_{1},a]\hookrightarrow\mathrm{Im}[x_{2},a]\xrightarrow{}Z\to 0$
where $Z=\mathrm{Im}[x_{2},a]/\mathrm{Im}[x_{1},a]\subseteq f_{a}/\mathrm{Im}[x_{1},a]$. The quotient $f_{a}/\mathrm{Im}[x_{1},a]$ is the cokernel of $f(x_{1}\leq a)$, so it is free by hypothesis. This implies that $Z$ is also free, so the sequence splits, which yields $\mathrm{Im}[x_{2},a]=\mathrm{Im}[x_{1},a]+A$ for some complement $A$. Therefore, if there exists infinitely many distinct submodules of form $\mathrm{Im}[x,a]$, the module $f_{a}$ would have infinite rank — a contradiction. Thus, the family $\{\mathrm{Im}[x,a]\}_{x\in I}$ of images into $f_{a}$ is finite. ∎
2.2. A brief discussion on finitely-indexed persistence modules
We briefly review relevant results on interval decompositions of finitely-indexed persistence modules with PID coefficients. See [10] for a more thorough discussion.
Consider the finite totally-ordered poset category $\mathcal{I}=\{0,\ldots,n\}$ with the usual ordering, and let $f:\mathcal{I}\to R\mathrm{-Mod}$ be a persistence module that is pointwise free and finitely-generated, where $R$ is a PID.
If the structure maps of $f$ all have free cokernels, then [10, Theorem 17] states that, for each pair $(p,q)$ of integers such that $1\leq p,q\leq n$, there exits a (non-unique) submodule $A^{pq}_{a}\subseteq f_{a}$ such that
$\mathrm{IK}_{a}[p,q]=$ $A^{pq}_{a}\oplus\Bigl(\mathrm{IK}_{a}[p-1,q]+\mathrm{IK}_{a}[p,q-1]\Bigr)$
Theorem [10, Theorem 18] states that, independent of one’s choice of $A^{pq}_{a}$, one has
(2.1) $\bigoplus_{p,q}A_{a}^{pq}$ $=f_{a}\,,\quad\bigoplus_{q\leq y}A_{a}^{pq}=\mathrm{Ker}[a,y]\,,\quad\bigoplus_{p\leq x}A_{a}^{pq}=\mathrm{Im}[x,a]\,$,
where $p,q\in\{1,\ldots,n\}$.
In the context of persistent homology, a submodule $A_{a}^{pq}\subseteq f_{a}$ in this decomposition can be thought of as cycles that are born at $p$ and die at $q$. This is because (1) whenever $a\notin[p,q)$, we must have $A_{a}^{pq}=0$ (see [10, Lemma 15]), and (2) one can choose submodules $A^{pq}_{a}$ such that $f(a\leq b)(A_{a}^{pq})=A_{b}^{pq}$. That is, for any choice of $A_{a}^{pq}$, the image $f(a\leq b)(A_{a}^{pq})$ is a complement of $\mathrm{IK}_{b}[p-1,q]+\mathrm{IK}_{b}[p,q-1]$ in $\mathrm{IK}_{b}[p,q]$ (see [10, Lemma 16]).
3. Necessity
Here we prove the necessity direction of Theorem 1.1.
Theorem 3.1.
Let $f:I\to R\mathrm{-Mod}$ be a persistence module that is pointwise free and finitely-generated, where $R$ is a PID. If $f$ admits an interval decomposition, then every structure map $f(a\leq b)$ has free cokernel.
Proof.
It is readily checked that for any persistence modules $h_{1},h_{2}:I\to R\mathrm{-Mod}$, we have $\text{coker}\Bigl((h_{1}\oplus h_{2})(a\leq b)\Bigr)\cong\mathrm{coker}\Bigl(h_{1}(a\leq b)\Bigr)\oplus\mathrm{coker}\Bigl(h_{2}(a\leq b)\Bigr)$. The cokernel of every structure map of an interval module is free, as the image is either $R$ or 0. Thus, the cokernel of every direct sum of interval modules is also free. ∎
4. Sufficiency
We now prove sufficiency of Theorem 1.1, which is the following.
Theorem 4.1.
Let $R$ be a PID and $f:I\to R\mathrm{-Mod}$ be a persistence module that is pointwise free and finitely-generated. Suppose every structure map $f(a\leq b)$ has free cokernel. Then $f$ admits an interval decomposition.
Proof.
This result depends on observations developed below, synthesized here into a complete proof. Suppose that the persistence module $f$ has free cokernels. Lemma 4.3 implies that $f$ admits a consistent basis (see Definition 4.2) if, for every proper subset $J\subsetneq I$ such that $\{-\infty,\infty\}\subseteq J$ and every consistent basis $(\beta_{j})_{j\in J}$ with respect to $f$, the basis can be extended by one element $\beta_{a}$, for $a\notin J$. This condition holds by Theorem 4.6. Therefore a consistent basis exists. ∎
Throughout this section, we fix a PID $R$, our coefficent ring; fix a persistence module $f:I\to R\mathrm{-Mod}$ that is pointwise free and finitely-generated; assume that every structure map has free cokernel.
4.1. Consistent bases and Zorn’s condition
We begin by establishing some common language to talk about bases for the modules of interest.
Definition 4.2.
  • The restriction of $f$ to $J$ is the restriction of the functor $f$ to the full subcategory $J\subseteq I$. We denote this object by $f|_{J}$.
  • A consistent basis for $f$ is an indexed family $(\beta_{i})_{i\in I}$ of bases such that the matrix representation of every structure map with respect to these bases is a matching matrix. (Note that $f$ has a consistent basis if and only if $f$ admits an interval decomposition; we use these terminologies interchangeably.)
  • A consistent basis of $f|_{J}$ with respect to $f$ is a consistent basis $(\beta_{j})_{j\in J}$ for $f|_{J}$, with the additional condition that, for each $j\in J$, the basis $\beta_{j}$ contains a basis for the submodules $\mathrm{Im}[i,j]$ and $\mathrm{Ker}[j,k]$, for all $i,k\in{I}$.
We address Theorem 4.1 by examining restrictions of $f$ and their consistent bases with respect to $f$.
Lemma 4.3 (Zorn’s condition).
The persistence module $f$ has a consistent basis if the following condition holds: For any set $J$ such that (i) $\{-\infty,\infty\}\subseteq J\subsetneq I$ and (ii) $f|_{J}$ admits a consistent basis $(\beta_{j})_{j\in{J}}$ with respect to $f$, there exists an element $a\in{I}\setminus{J}$ and a basis $\beta_{a}$ for $f_{a}$ such that “appending” $\beta_{a}$ to $(\beta_{i})_{i\in J}$ produces a consistent basis $(\beta_{i})_{i\in{J}\cup\{a\}}$ of $f|_{{J}\cup\{a\}}$ with respect to $f$.
Proof.
We apply Zorn’s lemma. Let $P$ be the set whose elements are pairs $({K},(\beta_{k})_{k\in{K}})$ such that $\{-\infty,\infty\}\subseteq{K}\subseteq{I}$ and $(\beta_{k})_{k\in{K}}$ is a consistent basis of $f|_{K}$ with respect to $f$. Impose a partial order on $P$ such that $({K},(\beta_{k})_{k\in{K}})\leq({K}^{\prime},(\beta^{\prime}_{k})_{k\in{K}^{\prime}})$ if and only if ${K}\subseteq{K}^{\prime}$ and $\beta_{k}=\beta^{\prime}_{k}$ for all $k\in{K}$. Note that $P$ is nonempty, as we can take $K=\{-\infty,\infty\}$ with $\beta_{-\infty}=\beta_{\infty}=\emptyset$ (recall that $f_{-\infty}=f_{\infty}=0$ by convention). Every chain in this poset clearly has an upper bound in $P$, as we can take the union of indexing sets. Therefore, by Zorn’s lemma, $P$ contains a maximal element.
Suppose, now, that the condition holds. Then $({K},(\beta_{k})_{k\in{K}})$ can only be maximal in $P$ if K=I{K}={I}, because if ${K}\subsetneq{I}$, we can find $a\in{I}\setminus{K}$ and a basis $\beta_{a}\subseteq f_{a}$ such that $({K}\cup\{a\},(\beta_{k})_{k\in{K}\cup\{a\}})\in P$. Therefore $f$ has a consistent basis, as desired. ∎
4.2. Partitioning consistent bases with images and kernels
We now partition a consistent basis $\beta_{j}$ into disjoint subsets $\beta_{j}^{pq}$, where each subset contains basis elements with the same “birth values” and “death values”.
Lemma 4.4.
Let $M$ be a free and finitely-generated $R$-module, and let $\beta\subseteq M$ be a basis. Let $\mathcal{M}_{\beta}$ be the family of submodules of $M$ that can be written as the span of a subset of $\beta$. That is $\mathcal{M}_{\beta}=\{\mathrm{span}(\gamma):\gamma\subseteq\beta\}$. The family $\mathcal{M}_{\beta}$ is closed under sum and intersection.
Proof.
Without loss of generality, we can assume $M=R^{n}$, and $\beta$ consists of standard unit vectors. Let $N_{1},N_{2}\in\mathcal{M}_{\beta}$, in which $N_{1}=\mathrm{span}(\gamma_{1})$ and $N_{2}=\mathrm{span}(\gamma_{2})$. It is readily checked that $N_{1}+N_{2}=\mathrm{span}(\gamma_{1}+\gamma_{2})$ and $N_{1}\cap N_{2}=\mathrm{span}(\gamma_{1}\cap\gamma_{2})$. ∎
Lemma 4.4 implies the following property for consistent bases.
Corollary 4.5.
If a basis $\beta_{j}$ for $f_{j}$ contains a basis for every image into and kernel out of $f_{j}$, then $\beta_{j}$ also contains a basis for any submodule of $f_{j}$ that can be obtained by taking sums and intersections of submodules of form $\mathrm{Im}[i,j]$ and $\mathrm{Ker}[j,k]$.
Fix $J\subsetneq I$ and $a\in I\setminus J$. Choose a finite index set $S=\{i_{1},\ldots,i_{r}\}\subseteq I$ such that, for every $c\in I$, we have $\mathrm{Ker}[a,c]=\mathrm{Ker}[a,i_{t}]$ for some $t$ and $\mathrm{Im}[c,a]=\mathrm{Im}[i_{s},a]$ for some $s$. Such a finite set exist by Lemma 2.9. We further require $\{i_{1},\ldots,i_{r}\}$ to (1) contain $a$, (2) be of minimal size, and (3) contain as many elements of $J$ as possible. (That is, for any $j\in J$ such that $j\geq a$, there must be some $i_{q}\in J$ such that $\mathrm{Ker}[a,j]=\mathrm{Ker}[a,i_{q}]$. Similarly, for any $j\in J$ such that $j\leq a$, there must be some $i_{p}\in J$ such that $\mathrm{Im}[j,a]=\mathrm{Im}[i_{p},a]$.)
Supposing that one exists, fix a consistent basis $(\beta_{j})_{j\in J}$ for $f|_{J}$ with respect to $f$. For each $j\in J$, define functions $\mathcal{K}:f_{j}\to\mathbb{Z}$ by $\mathcal{K}(v)=\min\{q:v\in\mathrm{Ker}[j,i_{q}]\}$ and $\mathcal{I}:f_{j}\to\mathbb{Z}$ by $\mathcal{I}(v)=\min\{p:v\in\mathrm{Im}[i_{p},j]\}$. Then, for all $j\in J$ and all $(p,q)\in\{1,\ldots,r\}\times\{1,\ldots,r\}$, define
$\beta^{pq}_{j}:=\{v\in\beta_{j}:(\mathcal{I}(v),\mathcal{K}(v))=(p,q)\}\,$, $A^{pq}_{j}:=\textrm{span}(\beta^{pq}_{j})\,.$
Several observations are immediate, for any $j\in J$.
  • Basis $\beta_{j}$ partitions as a disjoint union $\beta_{j}=\bigsqcup_{p,q}\beta_{j}^{pq}$.
  • Consequently, $f_{j}=\bigoplus_{p,q}A^{pq}_{j}$.
  • For each $i_{n}$ the set $\bigcup_{q\leq n}\beta^{pq}_{j}=\{v\in\beta_{j}:\mathcal{K}({v})\leq n\}=\beta_{j}\cap\mathrm{Ker}[j,i_{n}]$ forms a basis for $\mathrm{Ker}[j,i_{n}]$. Recall that $\beta_{j}$ contains a basis for $\mathrm{Ker}[j,i_{n}]$ by hypothesis (see Definition 4.2).
  • Similarly, for each $i_{n}$, the set $\bigcup_{p\leq n}\beta^{pq}_{j}=\{v\in\beta_{j}:\mathcal{I}({v})\leq n\}=\beta_{j}\cap\mathrm{Im}[i_{n},j]$ forms a basis for $\mathrm{Im}[i_{n},j]$. Recall that $\beta_{j}$ contains a basis for $\mathrm{Im}[i_{n},j]$ by hypothesis (see Definition 4.2).
  • If $V\subseteq f_{j}$ is any submodule which can be obtained via a finite sequence of sums and intersections of submodules of the form $\mathrm{Im}[i_{m},j]$ and $\mathrm{Ker}[j,i_{n}]$, then $V\cap\beta_{j}$ is a disjoint union of subsets of form ${\beta}^{pq}_{j}$. Moreover, $V\cap\beta_{j}$ is a basis for $V$. This observation follows from Lemma 4.4 and Corollary 4.5.
  • If $X$ and $Y$ are disjoint subsets of $\{1,\ldots,r\}\times\{1,\ldots,r\}$, then $\textrm{span}(\bigsqcup_{(p,q)\in X\cup Y}\beta^{pq}_{j})=\textrm{span}(\bigsqcup_{(p,q)\in X}\beta^{pq}_{j})\oplus\textrm{span}(\bigsqcup_{(p,q)\in Y}\beta^{pq}_{j})$
  • As a special case, we have
    $\mathrm{IK}_{j}[i_{p},i_{q}]$ $=\textrm{span}\left(\bigcup_{m\leq p,n\leq q}\beta_{j}^{pq}\right)$
    $=\textrm{span}(\beta^{pq}_{j})\oplus\textrm{span}\Biggl(\left(\cup_{m<p,n\leq q}\beta_{j}^{pq}\right)\bigcup\left(\cup_{m\leq p,n<q}\beta_{j}^{pq}\right)\Biggr)$
    $=\textrm{span}(\beta^{pq}_{j})\oplus\Biggl(\textrm{span}\left(\cup_{m<p,n\leq q}\beta_{j}^{pq}\right)+\textrm{span}\left(\cup_{m\leq p,n<q}\beta_{j}^{pq}\right)\Biggr)$
    (4.1) $=A^{pq}_{j}\oplus\Bigl(\mathrm{IK}_{j}[i_{p-1},i_{q}]+\mathrm{IK}_{j}[i_{p},i_{q-1}]\Bigr)$
4.3. Incremental Extension
Lemma 4.3 implies that, to show Theorem 4.1, it is enough to prove the following.
Theorem 4.6 (Incremental Extension).
Let $J\subsetneq I$ be a proper subset containing -$\infty$ and $\infty$, and suppose $f|_{J}$ admits a consistent basis $(\beta_{j})_{j\in J}$ with respect to $f$. Then, for each $a\in I\setminus J$, there exists a basis $\beta_{a}\subseteq f_{a}$ such that $(\beta_{j})_{j\in J\cup\{a\}}$ forms a consistent basis for $f|_{J\cup\{a\}}$ with respect to $f$.
Proof.
In the discussion below we construct a basis $\beta_{a}$, and show that $(\beta_{j})_{j\in J\cup\{a\}}$ forms a consistent basis for $f|_{J\cup\{a\}}$ with respect to $f$ (Theorem 4.13). ∎
The remainder of this section is devoted to the details of the proof of Theorem 4.6. We will first construct the desired basis $\beta_{a}\subseteq f_{a}$ by patching together images and inverse images of subsets of the bases $\{\beta_{j}\}_{j\in J}$, which already exist over $J$. We will then show that $\beta_{a}$ has the required properties.
Fix a subset $J$ such that $\{-\infty,\infty\}\subseteq J\subsetneq I$, and choose some $a\notin J$. Choose a (finite) index set $S=\{i_{1},\ldots,i_{r}\}$ as directed in Section 4.2. Additionally, choose indices $s<k<t$ such that (i) $i_{t}\in J\cap S$ is the smallest index such that $i_{t}>a$, (ii) $i_{s}\in J\cap S$ is the largest index such that $i_{s}<a$, and (iii) $i_{k}=a$. Note that $i_{s}$ and $i_{t}$ must exist because $J$ contains -$\infty$ and $\infty$.
Remark 4.7.
Because $S$ is finite, the results from [10] apply to our analysis of $f|_{S}$.
Partition $\beta_{j}$ into disjoint subsets $\beta_{j}^{pq}\subseteq\beta_{j}$, and define $A_{j}^{pq}=\textrm{span}(\beta^{pq}_{j})\subseteq f_{j}$ as directed in Section 4.2, for each $j\in J$. Intuitively, $\beta_{j}^{pq}$ represents a subset of $\beta_{j}$ consisting of basis vectors with the same “birth value” (represented by the integer $p$) and “death value” (represented by the integer $q$).
Now define corresponding subsets $\beta^{pq}_{a}\subseteq f_{a}$ as follows. See Figure 1 for reference.
  1. (1)
    If $i_{p}>a$ or $i_{q}\leq a$, we define $\beta^{pq}$ to be empty. This makes intuitive sense, because $p$ and $q$ represent “birth values” and “death values”, respectively; no vectors in $f_{a}$ are born later than $a$ or die before $a$.
  2. (2)
    For $i_{p}\leq i_{s}$ we will define $\beta^{pq}_{a}=f(i_{s}\leq a)(\beta^{pq}_{i_{s}})$. That is, if the “birth value” $i_{p}$ occurs before $i_{s}$, then we obtain a set of basis vectors $\beta^{pq}_{a}$ by pushing forward some basis vectors in $\beta_{i_{s}}$.
The remaining case is $i_{s}<i_{p}\leq a$. We divide this into the two subcases where $i_{t}<i_{q}$ and $i_{t}\geq i_{q}$, as follows:
  1. (3)
    For $i_{s}<i_{p}\leq a<i_{t}<i_{q}$, we define $\beta^{pq}_{a}\subseteq f(a\leq i_{t})^{-1}(\beta^{pq}_{i_{t}})$ by choosing one representative from the preimage of each basis element $v\in\beta^{pq}_{i_{s}}$. That is, if the “death value” $i_{q}$ occurs strictly after $i_{t}$, then we obtain a set of basis vectors $\beta^{pq}_{a}$ by pulling back some basis vectors in $\beta_{i_{t}}$. We require each preimage vector to lie in the submodule $\mathrm{IK}_{a}[i_{p},i_{q}]$. Note that this requirement can always be satisfied because each $v\in\beta^{pq}_{i_{t}}$ lies in the image of $f(i_{p}\leq i_{t})$, hence also in the image of $f(a\leq i_{t})$.
  2. (4)
    If none of the preceding rules applies (i.e., $i_{s}<i_{p}\leq a<i_{q}\leq i_{t}$), we choose any complement $A_{a}^{pq}$ of $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}[i_{p},i_{q-1}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$, and define $\beta^{pq}_{a}$ to be any basis of $A_{a}^{pq}$. (Such a complement exists by [10, Theorem 17].) In this case, there is nothing in $J$ to push forward or pull back to index $a$, so we choose an arbitrary complement.
Lemma 4.8.
The set $\beta^{pq}_{a}$ is linearly independent, for all $p$ and $q$.
Proof.
In case (1), $\beta^{pq}_{a}$ is empty, so there is nothing to prove.
In case (2), we have, by [10, Lemma 16], that $f(i_{s}\leq a)|_{{A_{i_{s}}^{pq}}}$ restricts to an isomorphism onto its image. Because $\beta^{pq}_{i_{s}}$ is linearly independent (it is a basis of $A^{pq}_{i_{s}}$), its image $f(i_{s}\leq a)(\beta^{pq}_{i_{s}})$ is as well.
In case (3), the elements of $\beta^{pq}_{a}$ map bijectively onto a linearly independent set $\beta^{pq}_{i_{t}}$, under a linear map; therefore, $\beta^{pq}_{a}$ is linearly independent.
In case (4), the set $\beta^{pq}_{a}$ is linearly independent by hypothesis. ∎
Lemma 4.9.
Let $A^{pq}_{a}:=\mathrm{span}(\beta^{pq}_{a})$. Then, for all $1\leq p,q\leq n$ we have
$\mathrm{IK}_{a}[i_{p},i_{q}]=A^{pq}_{a}\oplus\Biggl(\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]\Biggr)\,.$
Proof.
We check the four cases defined in the construction of $\beta^{pq}_{a}$.
Case (1) Suppose that $i_{p}>a$ or $i_{q}\leq a$. By the proof of [10, Lemma 15], we have that $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]=\mathrm{IK}_{a}[i_{p},i_{q}]$. This implies that the complement is 0, with basis $\beta^{pq}_{a}:=\emptyset$, as desired.
Case (2) Suppose that $p\leq s$. By [10, Lemma 16], $f(i_{s}\leq a)(A_{i_{s}}^{pq})$ is a complement of $\mathrm{IK}_{a}[i_{p},i_{q-1}]+\mathrm{IK}_{a}[i_{p-1},i_{q}]$ in $\mathrm{IK}_{a}[i_{p},i_{q}]$.
Case (3) Suppose that $i_{s}<i_{p}\leq a<i_{t}<i_{q}$. We first recall from (4.1) that $\mathrm{IK}_{i_{t}}[i_{p},i_{q}]=A^{pq}_{i_{t}}\oplus\Bigl(\mathrm{IK}_{i_{t}}[i_{p-1},i_{q}]+\mathrm{IK}_{i_{t}}[i_{p},i_{q-1}]\Bigr)$. Note that $A_{a}^{pq}$ can be realized as $\ell(A^{pq}_{{i_{t}}})$, where $\ell$ is some choice of split map in the short exact sequence in Lemma A.1. Therefore, by Lemma A.1, we have
$\mathrm{IK}_{a}[i_{p},i_{q}]$ $=\ell(A_{{i_{t}}}^{pq})\oplus\Bigl(\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]\Bigr)$
$=A_{a}^{pq}\oplus\Bigl(\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]\Bigr)$
Case (4) This case holds by definition.
As all cases have been addressed, we conclude the proof. ∎
Figure 1. The four cases for defining $\beta_{a}^{pq}$, for different values of $i_{p}$ and $i_{q}$.
Lemma 4.10.
The set $\beta_{a}:=\bigsqcup_{1\leq p,q\leq r}\beta^{pq}_{a}$ forms a basis for $f_{a}$.
Proof.
To show that $\beta_{a}=\bigsqcup_{p,q}\beta^{pq}_{a}$ is a basis for $f_{a}$, it is enough to show that $f_{a}=\bigoplus_{p,q}A^{pq}_{a}$. This is given by [10, Theorem 18]. ∎
Lemma 4.11.
The matrix representation $f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}$ with respect to bases $\beta_{i_{s}}$ and $\beta_{a}$ is a matching matrix. Similarly, the matrix representation $f(a\leq i_{t})]_{\beta_{a}}^{\beta_{i_{t}}}$ with respect to bases $\beta_{a}$ and $\beta_{i_{t}}$ is a matching matrix.
Proof.
We first show that $f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}$ is a matching matrix.
For each $p$ and $q$, one of the following mutually exclusive cases must hold:
  1. (1)
    $i_{p}>i_{s}$ or $i_{q}\leq i_{s}$. In this case, the basis $\beta^{pq}_{i_{s}}$ is empty, by [10, Lemma 15].
  2. (2)
    $i_{p}\leq i_{s}<i_{q}$. We divide this into two subcases: $a<i_{q}$ and $a\geq i_{q}(>i_{s})$
    1. (a)
      $a<i_{q}$. In this case, the structure map $f(i_{s}\leq a)$ maps $A^{pq}_{i_{s}}$ isomorphically onto $A^{pq}_{a}$ and, in particular, maps $\beta^{pq}_{i_{s}}$ bijectively onto $\beta_{a}^{pa}$.
    2. (b)
      $i_{s}<i_{q}\leq a$. In this case, $A^{pq}_{i_{s}}\subseteq\mathrm{Ker}[i_{s},a]$, so the corresponding basis $\beta_{i_{t}}^{pq}$ maps to 0.
Thus, in each case, the basis $\beta^{pq}_{i_{s}}$ is empty, maps to zero, or maps bijectively onto $\beta^{pq}_{a}$. Because $\beta_{a}=\bigsqcup_{p,q}\beta^{pq}_{a}$ and $\beta_{i_{s}}=\bigsqcup_{p,q}\beta^{pq}_{i_{s}}$, it follows that the map $f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}$ is a matching matrix.
Next, we show that $f(a\leq i_{t})]_{\beta_{a}}^{\beta_{i_{t}}}$ is a matching matrix. For each $p$ and $q$, one of the following mutually exclusive cases must hold.
  1. (1)
    $i_{p}>a$ or $i_{q}\leq a$. In this case, $A^{pq}_{a}=0$ by definition, with basis $\beta^{pq}_{a}=\emptyset$.
  2. (2)
    $i_{q}>i_{t}(>a)$. We divide this case into the subcases where $i_{s}\geq i_{p}$ and where $i_{s}<i_{p}(\leq a)$.
    1. (a)
      $i_{s}\geq i_{p}$. In this case, recall that $A_{a}^{pq}=f(i_{s}\leq a)(A^{pq}_{i_{s}})$, with basis $\beta^{pq}_{a}=f(a\leq b)(\beta^{pq}_{i_{s}})$, by construction. Because $(\beta_{j})_{j\in J}$ is a consistent basis with respect to $f$, we have $f(i_{s}\leq i_{t})$ mapping $\beta_{i_{s}}^{pq}$ bijectively onto $\beta_{i_{t}}^{pq}$. Therefore, functoriality of $f$ implies that $f(a\leq i_{t})$ maps $\beta_{a}^{pq}$ bijectively onto $\beta_{i_{t}}^{pq}$.
    2. (b)
      $i_{s}<i_{p}(\leq a)$. In this case, recall that we constructed the basis $\beta_{a}^{pq}\subseteq A_{a}^{pq}$ by taking the basis $\beta_{i_{t}}^{pq}\subseteq A_{i_{t}}^{pq}$ and choosing a preimage for each element. This implies that $f(a\leq i_{t})$ maps $\beta_{a}^{pq}$ bijectively into $\beta_{i_{t}}^{pq}$.
  3. (3)
    $i_{t}\geq i_{q}>a$. Note that $A^{pq}_{a}\subseteq\mathrm{Ker}[a,i_{t}]$. Therefore, $f(a\leq i_{t})$ maps the corresponding basis $\beta_{a}^{pq}$ to 0.
Thus, in each case, the basis $\beta^{pq}_{a}$ is empty, maps to zero, or maps bijectively onto $\beta^{pq}_{i_{t}}$. Because $\beta_{a}=\bigsqcup_{p,q}\beta^{pq}_{a}$ and $\beta_{i_{t}}=\bigsqcup_{p,q}\beta^{pq}_{i_{t}}$, it follows that the map $f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}$ is a matching matrix. ∎
We can extend Lemma 4.11 by the following corollary.
Corollary 4.12 (Matching matrices for $(S\cap J)\cup\{a\}$).
For any $i_{p}\in S\cap J$ such that $i_{p}<a$, the matrix representation $f(i_{p}\leq a)(A)]_{\beta_{i_{p}}}^{\beta_{a}}$ with respect to bases $\beta_{i_{p}}$ and $\beta_{a}$ is a matching matrix. Similarly, for any $i_{q}\in S\cap J$ such that $i_{q}>a$, the matrix representation $f(a\leq i_{q})(A)]_{\beta_{a}}^{\beta_{i_{q}}}$ with respect to bases $\beta_{a}$ and $\beta_{i_{q}}$ is a matching matrix.
Proof.
For any $i_{p}\in S\cap J$ such that $i_{p}<a$, observe that $f(i_{p}\leq a)=f(i_{s}\leq a)\circ f(i_{p}\leq i_{s})$. So, $f(i_{p}\leq a)]_{\beta_{i_{p}}}^{\beta_{a}}=[f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}[f(i_{p}\leq i_{s})]_{\beta_{i_{p}}}^{\beta_{i_{s}}}$. By Lemma 4.11, the matrix representation $f(i_{s}\leq a)]_{\beta_{i_{s}}}^{\beta_{a}}$ is a matching matrix, and $f(i_{p}\leq i_{s})]_{\beta_{i_{p}}}^{\beta_{i_{s}}}$ is a matching matrix by assumption that $\{\beta_{j}\}_{j\in J}$ is a consistent basis with respect to $f$. Therefore, it follows that $f(i_{p}\leq a)]_{\beta_{i_{p}}}^{\beta_{a}}$ is a matching matrix.
Similarly, for any $i_{q}\in S\cap J$ such that $i_{q}>a$, observe that $f(a\leq i_{q})=f(i_{t}\leq i_{q})\circ f(a\leq i_{t})$. By a similar argument, $f(a\leq i_{q})]_{\beta_{a}}^{\beta_{i_{q}}}$ is a matching matrix. ∎
We now show $(\beta_{j})_{j\in J\cup\{a\}}$ forms a consistent basis for $f|_{J\cup\{a\}}$ with respect to $f$.
Theorem 4.13.
The family of bases $(\beta_{j})_{j\in J\cup\{a\}}$ forms a consistent basis for $f|_{J\cup\{a\}}$ with respect to $f$.
We show Theorem 4.13 in two steps. We first show, in Theorem 4.14, that the matrix representation of every structure map with respect to the bases $(\beta_{j})_{j\in J\cup\{a\}}$ is a matching matrix. Next, we show, in Theorem 4.16, that every $\beta_{j}$, where $j\in J\cup\{a\}$, contains a basis for all of the kernel and image submodules that are required for a basis to meet the criteria of a consistent basis with respect to $f$.
Figure 2. Indices in the proof of Theorem 4.14. For $c,b\in J$ such that $c<a<b$, we choose $i_{u},i_{\ell}\in S$ such that $\mathrm{Im}[c,a]=\mathrm{Im}[i_{u},a]$ and $\ker[a,b]=\ker[a,i_{\ell}]$. The index $i_{u}$ (respectively, $i_{\ell}$) may be greater or less than $c$ (respectively, $b$).
Theorem 4.14 (Matching matrices for $J\cup\{a\}$).
The matrix representation of every structure map with respect to the bases $(\beta_{j})_{j\in J\cup\{a\}}$ is a matching matrix. That is, for every $c,b\in J\cup\{a\}$, the matrix representation $f(c\leq b)]_{\beta_{c}}^{\beta_{b}}$ is a matching matrix.
Proof.
This proof involves several elements of $I$, labeled $a,b,c,i_{u},$ and $i_{\ell}$. See Figure 2 for a visual aid of the relative positions of these indices.
It is enough to focus only on structure maps that start or end at $f_{a}$, because $(\beta_{j})_{j\in J}$ (i.e., without $a$) is assumed to be a consistent basis with respect to $f$.
We first look at maps that end at $f_{a}$. Take any $c\in J$ such that $c<a$. Choose $i_{u}\in S\cap J$ such that $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. If $c\leq i_{u}\leq a$, then $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}=[f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}[f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$. Note that $f(c\leq i_{u})]_{\beta_{c}}^{\beta_{i_{u}}}$ is a matching matrix because $i_{u},c\in J$, and $(\beta_{j})_{j\in J}$ is a consistent basis. Additionally, by Corollary 4.12, $f(i_{u}\leq a)]_{\beta_{i_{u}}}^{\beta_{a}}$ is a matching matrix. Therefore $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Now, consider $i_{u}\leq c<a$, and let $\{v_{1},\ldots,v_{r}\}:=\beta_{i_{u}}\setminus\mathrm{Ker}[i_{u},a]$. By Corollary 4.12, this set maps injectively into $\beta_{a}$, and its image under $f(i_{u}\leq a)$ spans $\mathrm{Im}[i_{u},a]=\mathrm{Im}[c,a]$. By functoriality, this implies that the image of the set $\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}\subseteq\beta_{c}$ maps injectively into the basis $\beta_{a}\subseteq f_{a}$ under $f(c\leq a)$. Because $\beta_{c}$ contains a basis for $\mathrm{Ker}[c,a]$, we have, by the rank-nullity theorem, that $\beta_{c}\setminus\{f(i_{u}\leq c)(v_{1}),\ldots,f(i_{u}\leq c)(v_{r})\}$ spans $\mathrm{Ker}[c,a]$. Therefore, it follows that $f(c\leq a)]_{\beta_{c}}^{\beta_{a}}$ is a matching matrix.
Next, we look at maps that begin at $f_{a}$. Take any $b\in J$ such that $b>a$. By definition, there exists $i_{\ell}\in S\cap J$ such that $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$. If $a<i_{\ell}\leq b$, then $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}=[f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}[f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ is a matching matrix because both $f(i_{\ell}\leq b)]_{\beta_{i_{\ell}}}^{\beta_{b}}$ and $f(a\leq i_{\ell})]_{\beta_{a}}^{\beta_{i_{\ell}}}$ (by Corollary 4.12) are matching matrices.
Now, suppose that $a<b\leq i_{\ell}$. Let $v\in\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]$ be given. Because $\mathrm{Ker}[a,i_{\ell}]=\mathrm{Ker}[a,b]$, the elements $f(a\leq b)(v)$ and $f(a\leq i_{\ell})(v)$} are nonzero. We know that $f(a\leq i_{\ell})(v)\in\beta_{i_{\ell}}$, by Corollary 4.12. Therefore, because $(\beta_{j})_{j\in J}$ is a consistent basis, there exists $w\in\beta_{b}$ such that $f(b\leq i_{\ell})(w)=f(a\leq i_{\ell})(v)$.
We will show that $f(a\leq b)(v)=w$. To see this, note that because $\mathrm{Ker}[a,b]=\mathrm{Ker}[a,i_{\ell}]$, functoriality implies that $\mathrm{Im}[a,b]\cap\mathrm{Ker}[b,i_{\ell}]=0$. Let $\{w_{1},\ldots,w_{r}\}\subseteq\beta_{b}$ be a basis for $\mathrm{Im}[a,b]$, in which case, we can write $f(a\leq b)(v)=c_{1}w_{1}+\ldots+c_{r}w_{r}$. By applying $f(b\leq i_{\ell})$ to both sides, we get $f(a\leq i_{\ell})(v)=c_{1}f(b\leq i_{\ell})(w_{1})+\ldots+c_{r}f(b\leq i_{\ell})(w_{r})$. Because $\mathrm{Im}[a,b]\cap\mathrm{Ker}[b,i_{\ell}]=0$ and $(\beta_{j})_{j\in J}$ is a consistent basis with respect to $f$, we must have $f(b\leq i_{\ell})(w_{1}),\ldots,f(b\leq i_{\ell})(w_{r})\in\beta_{i_{\ell}}$, and because we also have $f(a\leq i_{\ell})(v)\in\beta_{i_{\ell}}$, it follows that all but one of the $c_{i}$ are 0, with the remaining one being $1$. Without loss of generality, suppose $c_{1}=1$ and $c_{2}=\ldots=c_{r}=0$, in which case, we have $f(a\leq i_{\ell})(v)=f(b\leq i_{\ell})(w_{1})$; because $w\in\beta_{b}$ is the unique basis element in $\beta_{b}$ that maps to $f(a\leq i_{\ell})(v)$, it follows that $w=w_{1}$. This yields $f(a\leq b)(v)=w$, as desired.
From this, it follows that $f(a\leq b)$ maps $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$ into $\beta_{b}$. Moreover, functoriality implies that this map is injective, because the composite map $f(a\leq i_{\ell})=f(b\leq i_{\ell})\circ f(a\leq b)$ is injective on $\beta_{a}\setminus\mathrm{Ker}[a,i_{\ell}]=\beta_{a}\setminus\mathrm{Ker}[a,b]$, by Corollary 4.12. This implies that $f(a\leq b)]_{\beta_{a}}^{\beta_{b}}$ is a matching matrix, as desired. ∎
Next, we show that for every $j\in J\cup\{a\}$, the basis $\beta_{j}$ contains a basis for every $\mathrm{Ker}[j,b]$ and every $\mathrm{Im}[c,j]$. We know that for $j\in J$, the basis $\beta_{j}$ contains a basis for every $\mathrm{Ker}[j,b]$ and every $\mathrm{Im}[c,j]$ because the indexed family $(\beta_{j})_{j\in J}$ (without the basis $\beta_{a}$ appended) is assumed to be a consistent basis with respect to $f$ (see Definition 4.2). Therefore, we only need to show this for $j=a$.
We use the following fact.
Lemma 4.15.
Suppose that $g:K\to R\mathrm{-Mod}$ is a persistence module with consistent basis $(\gamma_{k})_{k\in K}$. For any $a\in K$, the basis $\gamma_{a}$ contains a basis for every submodule of the form $\mathrm{Im}[c,a]$ and $\mathrm{Ker}[a,b]$.
Proof.
For any $x,y\in K$ such that $x\leq y$, define $\mathcal{K}_{x,y}=\gamma_{x}\cap\mathrm{Ker}[x,y]$, and $\mathcal{I}_{x,y}=\gamma_{x}\setminus\mathcal{K}_{x,y}$. Let $c\leq a\leq b$ be given. Clearly, $f(c\leq a)(\mathcal{I}_{c,a})$ spans the image submodule $\mathrm{Im}[c,a]$. Because $f(c\leq a]_{\gamma_{c}}^{\gamma_{a}}$ is a matching matrix, the set $\mathcal{I}_{c,a}$ maps injectively into $\gamma_{a}$, and therefore, $f(c\leq a)(\mathcal{I}_{c,a})\subseteq\gamma_{a}$ forms a basis for $\mathrm{Im}[c,a]$. Moreover, by the rank nullity theorem, $\mathcal{N}_{a,b}$ forms a basis for $\mathrm{Ker}[a,b]$. ∎
Theorem 4.16.
The basis $\beta_{a}$ for $f_{a}$ contains a basis for every $\mathrm{Ker}[a,b]$ and $\mathrm{Im}[c,a]$.
Proof.
Corollary 4.12 already established that $(\beta_{j})_{j\in S}$ is a consistent basis for $f|_{S}$. Therefore, by Lemma 4.15, the set $\beta_{a}$ contains a basis for $\mathrm{Im}[s,a]$ and $\mathrm{Ker}[a,s]$ for all $s\in S$. By hypothesis, every submodule of the form $\mathrm{Ker}[a,b]$ (respectively, $\mathrm{Im}[c,a]$) can be expressed in the form $\mathrm{Ker}[a,s]$ (respectively, $\mathrm{Im}[s,a]$) for some $s\in S$. The desired conclusion follows. ∎
5. Application: Integer decomposition, relative homology, and field-choice independence in persistent homology
We relate the decomposability of integer persistent homology to the sensitivity of persistence diagrams to choice of coefficient field. Let $\mathcal{K}$ be a filtration, viewed as a functor from $I$ to the category of topological spaces and inclusions. Let $\mathcal{H}_{n}(\mathcal{K};F)$ be the functor obtained by composing $\mathcal{K}$ with the $n$th-homology functor over a field $F$. The persistence diagram of $\mathcal{K}$ with coefficients in $F$, denoted $\mathrm{PD}_{n}^{F}(\mathcal{K})$, is the multiset of intervals corresponding to an interval decomposition of $\mathcal{H}_{n}(\mathcal{K};F)$. The field-independence problem asks: under what conditions is the persistence diagram $\mathrm{PD}_{n}^{F}(\mathcal{K})$ independent of the choice of $F$? Obayashi and Yoshiwaki [12] addressed this in the special case where $I$ is finite. We extend this to the infinite case.
Theorem 5.1.
Suppose that the homology groups $H_{n-1}(\mathcal{K}_{a};\mathbb{Z})$ and $H_{n}(\mathcal{K}_{a};\mathbb{Z})$ are free and finitely-generated for all $a\in I$. Given any $a\leq b$, consider the following commutative diagram, where ()(*) are extension of scalars and $(\dagger)$ arise from the universal coefficient theorem.
(5.1) $\mathcal{H}_{n}(\mathcal{K}_{a};\mathbb{Z})$$\mathcal{H}_{n}(\mathcal{K}_{a};\mathbb{Z})\otimes F$$\mathcal{H}_{n}(\mathcal{K}_{a};F)$$\mathcal{H}_{n}(\mathcal{K}_{b};\mathbb{Z})$$\mathcal{H}_{n}(\mathcal{K}_{b};\mathbb{Z})\otimes F$$\mathcal{H}_{n}(\mathcal{K}_{b};{F})$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$$\scriptstyle{(*)}$$\scriptstyle{(\dagger)}\scriptstyle{\cong}$
The maps $(\dagger)$ are isomorphisms, and the following are equivalent.
  1. (1)
    The module $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules.
  2. (2)
    The cokernel of each induced map $H_{n}(\mathcal{K}_{a};\mathbb{Z})\to H_{n}(\mathcal{K}_{b};\mathbb{Z})$ is free.
  3. (3)
    For all $a\leq b$ in $I$, the relative homology group $H_{n}(\mathcal{K}_{b},\mathcal{K}_{a};\mathbb{Z})$ is free.
  4. (4)
    The persistence diagram $\textrm{PD}_{n}^{F}(\mathcal{K})$ is identical for every coefficient field $F$.
Proof.
If $\textrm{PD}_{n}^{F}(\mathcal{K})$ is identical for every coefficient field $F$, then the same holds for $\mathcal{K}|_{S}$, for every finite subset $S\subseteq I$. It follows by [12, Theorem 1.9] that (4) implies (3). The argument from [10, Theorem 4] carries over without modification to show that (3) implies (2). Our main result, Theorem 1.1, shows that (2) implies (1).
Because $H_{n-1}(\mathcal{K}_{a};\mathbb{Z})$ and $H_{n}(\mathcal{K}_{a};\mathbb{Z})$ are free for all $a$, the maps $(\dagger)$ in (5.1) are isomorphisms, by the universal coefficient theorem. The extension maps ()(*) induce homomorphisms $\varphi:\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\to\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\otimes F$ of persistence modules, which preserves direct sums, in the sense that $\varphi(\bigoplus_{i}A_{i})=\bigoplus_{i}\varphi(A_{i})$. Therefore, the same holds for the map $\psi:\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\to\mathcal{H}_{n}(\mathcal{K};F)$ obtained by composing $\varphi$ with the isomorphism $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})\otimes F\xrightarrow{\cong}\mathcal{H}_{n}(\mathcal{K};F)$. Consequently, if $\mathcal{H}_{n}(\mathcal{K};\mathbb{Z})$ splits as a direct sum of interval submodules, then $\psi$ yields a corresponding decomposition of $\mathcal{H}_{n}(\mathcal{K};F)$, with the same multiset of intervals. This establishes that (1) implies (4). The same homomorphism carries consistent bases to consistent bases. ∎
Remark 5.2.
An analogous result holds for persistent cohomology.
Appendix A A useful short exact sequence
Here we present a short exact sequence used to prove Lemma 4.9. The notation is the same as in the proof of that result.
Lemma A.1.
Fix $i_{s}<i_{p}\leq a<i_{t}<i_{q}$, and consider the diagram
[…]
where $\iota$ is the inclusion map; $\pi$ is the projection map which restricts to zero on $\mathrm{IK}_{i_{t}}[i_{p-1},i_{q}]+\mathrm{IK}_{i_{t}}[i_{p},i_{q-1}]$ and to the identity map on $A^{pq}_{i_{t}}$; and $\phi=\pi\circ f(a\leq i_{t})|_{\mathrm{IK}_{a}[i_{p},i_{q}]}$. Then the sequence consisting of $\iota$ and $\phi$ is exact. Moreover, this sequence splits.
Proof.
The sequence is exact at $\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]$ because $\iota$ is injective.
To check exactness at $A_{i_{t}}^{pq}$, because $\pi$ is surjective, it suffices show that $f(a\leq i_{t})$ restricts to a surjection $\mathrm{IK}_{a}[i_{p},i_{q}]\to\mathrm{IK}_{i_{t}}[i_{p},i_{q}]$. By [10, Theorem 15], we have $f(a\leq i_{t})(\mathrm{IK}_{a}[i_{p},i_{q}])=\mathrm{IK}_{i_{t}}[i_{p},i_{q}]\cap\mathrm{Im}[a,i_{t}]$. However, because $i_{p}\leq a$, we have $\mathrm{Im}[i_{p},i_{t}]\subseteq\mathrm{Im}[a,i_{t}]$, which implies that
$\mathrm{IK}_{i_{t}}[i_{p},i_{q}]\cap\mathrm{Im}[a,i_{t}]$ $=\mathrm{Im}[i_{p},i_{t}]\cap\mathrm{Ker}[i_{t},i_{q}]\cap\mathrm{Im}[a,i_{t}]$
$=\mathrm{Im}[i_{p},i_{t}]\cap\mathrm{Ker}[i_{t},i_{q}]$
$=\mathrm{IK}_{i_{t}}[i_{p},i_{q}]\,{.}$
Finally, we show exactness at $\mathrm{IK}_{a}[i_{p},i_{q}]$. It is enough to show that
(A.1) $\Bigl(f(a\leq i_{t})|_{\mathrm{IK}_{a}[i_{p},i_{p}]}\Bigr)^{-1}\Bigl(\mathrm{IK}_{i_{t}}[i_{p-1},i_{q}]$ $+\mathrm{IK}_{i_{t}}[i_{p},i_{q-1}]\Bigr)$
$=\mathrm{IK}_{a}[i_{p-1},i_{q}]+\mathrm{IK}_{a}[i_{p},i_{q-1}]\,.$
Note that by [10, Theorem 16], we have
$f(a\leq{i_{t}})^{-1}\Bigl(\mathrm{IK}_{i_{t}}[i_{p-1},i_{q}]+\mathrm{IK}_{i_{t}}[i_{p},i_{q-1}]\Bigr)=\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]+\mathrm{Ker}[a,i_{t}]\,.$
Therefore, it follows that
$\Bigl(f(a\leq{i_{t}})|_{\mathrm{IK}_{a}[i_{p},i_{q}]}\Bigr)^{-1}$ $\Bigl(\mathrm{IK}_{i_{t}}[i_{p-1},i_{q}]+\mathrm{IK}_{i_{t}}[i_{p},i_{q-1}]\Bigr)$
(A.2) $=\Bigl(\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]+\mathrm{Ker}[a,i_{t}]\Bigr)\cap\mathrm{IK}_{a}[i_{p},i_{q}]$
(A.3) $=\Bigl(\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]\Bigr)+\Bigl(\mathrm{Ker}[a,i_{t}]\cap\mathrm{IK}_{a}[i_{p},i_{q}]\Bigr)$
(A.4) $=\Bigl(\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]\Bigr)+\mathrm{IK}_{a}[i_{p},i_{t}]$
(A.5) $=\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]\,.$
The equality between (A.2) and (A.3) follows from the modularity of submodules [^2] An order lattice is modular if, for any $a\leq b$, the identity $(a\vee x)\wedge b=a\vee(x\wedge b)$ holds for all $x$. The lattice of submodules of a module over a ring, ordered under inclusion, is always modular. In this case, we take $a=\mathrm{IK}_{{a}}[i_{p-1},i_{q}]+\mathrm{IK}_{{a}}[i_{p},i_{q-1}]$, $b=\mathrm{IK}_{{a}}[i_{p},i_{q}]$, and $x=\mathrm{Ker}[a,i_{{t}}]$. , and the equality between (A.3) and (A.4) holds because $i_{t}<i_{q}$ implies that $\mathrm{Ker}[a,i_{t}]\cap\mathrm{IK}_{a}[i_{p},i_{q}]=\mathrm{Ker}[a,i_{t}]\cap\mathrm{Ker}[a,i_{q}]\cap\mathrm{Im}[i_{p},a]=\mathrm{Ker}[a,i_{t}]\cap\mathrm{Im}[i_{p},a]=\mathrm{IK}_{a}[i_{p},i_{t}]$. The equality between (A.4) and (A.5) holds because $i_{t}<i_{q}$ implies that $\mathrm{IK}_{a}[i_{p},i_{t}]\subseteq\mathrm{IK}_{a}[i_{p},i_{q-1}]$. Therefore, our desired equality (A.1) follows.
This establishes that the sequence is exact. Moreover, the sequence splits because $A^{pq}_{i_{t}}$ is a submodule of a free module, and therefore is free. ∎